For the year ending 30 June 2019, Sandon Ltd reports net profit after tax of $2000 000. At the beginning of the year, Sandon Ltd had 1600 000 fully paid ordinary shares. It also had 200 000 $1.00, 20 per cent, cumulative preference shares outstanding. The preference shares were classified as equity.

On 1 September 2018 the company issued another 400 000 fully paid ordinary shares by way of a rights issue. The right provided an additional share for each four held, and required the payment of $3.00. The last rights share price was $4.

Required

Compute the basic EPS amount for 2019.
Answer #1

1. Ex-right Value = (Market Price *Number of shares) + (Exercise Price*Number of rights exercised) / Number of Shares after right issue

= (4*1,600,000) + (3*400,000) / 2,000,000 = 3.8

2. Bonus Factor = Fair Value per share immediately before right issue / Ex-right value per share = 4/3.8 = 1.05

3.

Date

Time proportion

(A)

Number of share

(B)

Bonus Factor

(C)

Adjusted number of shares

(D=B*C)

Weight average number of shares

(E=A*D)

1/07/2018 - 31/08/2018 2/12 1,600,000 1.05 1,680,000 280,000
1/09/2018 - 30/06/2019 10/12 2,000,000 1.00 2,000,000 1,666,667

4. Earnings available to equity shareholders = 2,000,000 - (200,000*20%) = 1,960,000

5. EPS = Earnings available to euity shareholders / Weighted average number of shares = 1,960,000/1,946,667

= $1 per share